OpenStudy (anonymous):

Sandra sets the cruise control in her car to 72 miles per hour when she goes on road trips. Which equation can be used to find how many hours, h, it will take her to drive m miles at her constant speed?

OpenStudy (anonymous):

72 m = 4 m/72=h 72+m=h 72/m=h

OpenStudy (gorv):

speed = distance/time

OpenStudy (gorv):

distance=m